# STAGE: Storyboard-Anchored Generation for Cinematic Multi-shot Narrative

### v1

@@ Abstract
While recent advancements in generative models have achieved remarkable visual fidelity in video synthesis, creating coherent multi-shot narratives remains a significant challenge. To address this, keyframe-based approaches have emerged as a promising alternative to computationally intensive end-to-end methods, offering the advantages of fine-grained control and greater efficiency. However, these methods often fail to maintain cross-shot consistency and capture cinematic language. In this paper, we introduce STAGE, a SToryboard-Anchored GEneration workflow to reformulate the keyframe-based multi-shot video generation task. Instead of using sparse keyframes, we propose STEP2 to predict a structural storyboard composed of start-end frame pairs for each shot. We introduce the multi-shot memory pack to ensure long-range entity consistency, the dual-encoding strategy for intra-shot coherence, and the two-stage training scheme to learn cinematic inter-shot transition. We also contribute the large-scale ConStoryBoard dataset, including high-quality movie clips with fine-grained annotations for story progression, cinematic attributes, and human preferences. Extensive experiments demonstrate that STAGE achieves superior performance in structured narrative control and cross-shot coherence.
Figure 1: Illustration of our proposed STAGE workflow for multi-shot video generation. Given a user-provided story description, existing end-to-end [cinetrans] and keyframe-based [iclora] methods often suffer from incoherent inter-shot transitions (e.g., motion discontinuities and appearance inconsistencies) that disrupt the narrative flow. We address this by explicitly predicting a structural storyboard composed of start-end frame pairs, which serve as visual anchors to ensure superior long-range shot consistency.
@@ 1 Introduction
Recent advancements in generative models have dramatically improved the visual quality of synthesized videos, fueling their adoption on short-clip content platforms [sora, sora2, kling]. To further unlock their potential for more sophisticated applications (e.g., filmmaking), recent models are evolving to enhance storytelling ability [storydiffusion, seedstory, synthesizing]. However, this narrative ability typically requires composing a long video from multiple distinct shots depicting different stages or perspectives of a story, which still poses significant challenges for existing models.
While directly generating an entire multi-shot video in an end-to-end manner [cinetrans, storyanchors, shotadapter, streamingt2v] is an intuitive approach, it is computationally expensive and runs in an “all-or-nothing” paradigm, leading to an inefficient trial-and-error process with limited user control. To provide fine-grained user control, keyframe-based approaches [cut2next, movieagent, videogenofthought, moviedreamer] first generate several keyframes to establish the video’s narrative structure, and subsequently employ external I2V models [dynamicrafter, gen4, svd] to synthesize each shot accordingly. However, these methods suffer from disrupted cross-shot coherence and struggle to capture cinematic language (Fig.˜1 (a), the close-up reveals motion discontinuity when opening the ring box).
@@
To address this challenge, we introduce STAGE, a workflow for SToryboard-Anchored GEneration for cinematic multi-shot narrative. The technical core of this workflow is our STart-End frame-Pair Prediction model (STEP2), which iteratively visualizes the text-based storyboard into $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$ pairs. To further enable STEP2 to visualize the storyboard of each shot effectively, we propose: (i) a multi-shot memory pack to compress the history and ensure long-range entity consistency; (ii) a dual-encoding strategy to enforce intra-shot coherence and logical correlation; and (iii) a two-stage training scheme to understand the complex cinematic language of inter-shot transitions. In the complete STAGE workflow, a director agent first augments the user-provided story theme into a text-based storyboard. Our fully-trained STEP2 model then iteratively generates the $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$ pairs. These pairs are finally fed into an off-the-shelf video generation model (e.g., WanX [wanx], Veo3.1 [veo3]) to generate video clips, which are then concatenated into the final multi-shot video.
To facilitate the training of STEP2, we collect open-sourced movie clips [condensedmovie] and construct the ConStoryBoard dataset [^1] We will publish the dataset once the paper is accepted.. It consists of 100K video clips with start-end frame pairs, along with structural storyboard annotations (e.g., story progress of each shot) and diverse cinematic language attributes (e.g., shot scale, shot length, camera angle, and camera movement). Building on this, we further curate the ConStoryBoard-HP by manually selecting high-quality frame pairs and constructing preference tuples for human-preference alignment post-training.
@@ 2.1 Single-Shot video generation
The field of single-shot video generation has seen remarkable progress. Building on the success of image diffusion models [stablediffusion, imagen, sdxl, pixart], early methods [imagenvideo, makeavideo, vdm] adapt pre-trained models by incorporating temporal layers. However, these approaches often result in limited temporal coherence. To address these limitations, research has shifted towards Diffusion Transformer architectures [scalable]. While state-of-the-art video foundation models [cogvideox, hunyuanvideo, wanx] are able to generate high-fidelity videos with a single shot, they typically lack a native mechanism for multi-shot narratives. Consequently, naively applying them shot-by-shot results in jarring visual discontinuities that disrupt the narrative flow, marking the key challenge our work addresses.
@@ 2.2 Multi-Shot video generation
Generating multi-shot videos requires maintaining both content consistency and narrative coherence across shots. Existing methods extend single-shot models to generate full sequences, either via end-to-end visual conditioning [videostudio, moviedreamer] or by modifying attention mechanisms [shotadapter, cinetrans, holocine]. These approaches are computationally expensive and offer limited user control. To address these limitations, an alternative approach [storydiffusion, seedstory, movieagent, cineverse] first generates sparse keyframes and then synthesizes the full video clips based on them. Despite their advances in text fidelity and story coherence, they largely neglect the cinematic language for inter-shot transitions, leading to abrupt cuts and logical disconnects that break the narrative flow. In contrast, our STAGE explicitly reformulates the keyframe-based approach as a start-end frame pair prediction problem, thereby directly modeling these inter-shot transitions.
@@ 2.3 Reinforcement learning for visual generation
Inspired by the success of Reinforcement Learning from Human Feedback (RLHF) in language models, aligning visual models with human preferences is a promising research area. Initial efforts involve directly fine-tuning models with scalar reward signals [imagereward, clark2023directly] or employing reward weighted regression [peng2019advantage, lee2023aligning]. Subsequently, policy gradient algorithms (e.g., PPO [ppo]) are integrated into diffusion models, demonstrating notable efficacy in visual quality improvement [fan2023reinforcement, miao2024training]. To avoid significant computational overhead and training instability, Direct Preference Optimization (DPO) [dpo] emerges and demonstrates its effectiveness in aligning visual models with human preferences for static image properties, such as text-image fidelity [wallace2024diffusion, wang2025diffusion] and aesthetic quality [liang2025aesthetic]. Building on this, we extend this technique to optimize for the complex temporal relationships (e.g., cinematic language) that define coherent storytelling in multi-shot videos.
@@ 3 Dataset
Current keyframe-based video datasets [cinetrans, storyboard, cineverse] offer a single keyframe for each shot, primarily focusing on text-to-shot alignment while neglecting the start and end frames for modeling inter-shot transitions. This motivates us to construct the ConStoryBoard dataset, tailored to train our start-end frame pair prediction model.
The construction of ConStoryBoard begins with multi-shot videos collected from the Condensed Movies dataset [condensedmovie], which we filter for high-resolution (over 1080p) and high aesthetic scores [laion] (over 5.5). These filtered videos are then segmented into individual video clips for each shot using TransNetV2 [transnetv2]. For each resulting clip ($i$-th shot in the original video), we employ InternVL-3.5 [internvl] to generate a detailed text description $D_{i}$ of the story progression, and structured cinematic attributes $C_{i}$ (e.g., shot scale and camera movement). We then extract the ground-truth start-end frame pair $(F^{\mathrm{S}}_{i},F^{\mathrm{E}}_{i})$ from these annotated clips, and post-process them by cropping black bars and removing watermarks. Finally, the ConStoryBoard dataset comprises 100K training pairs and 1K testing pairs, where each sample consists of a ground-truth pair $(F^{\mathrm{S}}_{i},F^{\mathrm{E}}_{i})$ and its corresponding annotations $(D_{i},C_{i})$.
@@ 4.1 Start-end frame pair prediction model
Specifically, as illustrated in Fig.˜3 (a), given preceding shot start-end frame pairs $\{(F_{j}^{\mathrm{S}},F_{j}^{\mathrm{E}})\}_{j=1}^{i-1}$ for $i$-th shot, we disorderly collect all frames in these pairs into a memory bank $\{F_{j}^{\mathrm{M}}\}_{j=1}^{2i-2}$, and encode all these frames into the latent space $\{m_{j}\}_{j=1}^{2i-2}=\{\mathcal{E}_{\mathrm{vae}}(F_{j}^{\mathrm{M}})\}_{j=1}^{2i-2}$ using a pre-trained VAE encoder $\mathcal{E}_{\mathrm{vae}}$. To guide the iterative generation with semantically relevant memory cues, we rank these preceding latent codes based on their CLIP similarity [clip] as $\{m^{\prime}_{j}\}_{j=1}^{2i-2}=\mathrm{Sort}\big(\{m_{j}\}_{j=1}^{2i-2}\big)$. After that, this ranked memory bank is further compressed into a packed memory token via a progressive spatial tiling mechanism:
@@
Specifically, as shown in Fig.˜3 (b), we encode the ground truth start-end frame pair $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$ separately using a pre-trained VAE encoder $\mathcal{E}_{vae}$, and then concatenate them along the sequence dimension to construct a single joint shot tensor $x_{i}=\big[\mathcal{E}_{\mathrm{vae}}(F_{i}^{\mathrm{S}});\mathcal{E}_{\mathrm{vae}}(F_{i}^{\mathrm{E}})\big]$. As the common practice of flow matching [lipman2022flow], this shot tensor is then linearly interpolated with a Gaussian noise as:
@@
Multi-modal understanding and generation. To reason about shot semantics from diverse context, our STEP2 first adopts a multi-modal large language model to understand the story progress and character performance of each shot. Specifically, as illustrated in Fig.˜3 (c), the understanding model $\mathcal{E}_{mu}$ is built upon Qwen2.5-VL [qwen2.5vl], which leverages the end frame of the previous shot $F_{i-1}^{\mathrm{E}}$, the $i$-th shot corresponding text description $D_{i}$, and cinematic attributes $C_{i}$ to generate unified context tokens $U_{i}=\mathcal{E}_{mu}(F_{i-1}^{\mathrm{E}},D_{i},C_{i})$ for the $i$-th shot.
With the aforementioned memory tokens $M_{i}$ and interpolated joint shot tensor $x_{i}^{t}$, these unified context tokens $U_{i}$ are fed into the generation model $\mathcal{E}_{gen}$ to generate the corresponding start-end frame pair $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$. Specifically, the generation model $\mathcal{E}_{gen}$ is implemented as a diffusion Transformer architecture, comprising multiple MMDiT blocks [esser2024scaling] that includes the self-attention layers for global context interaction. This enables our STEP2 to generate clean start-end frame pairs referring all integrated and compression semantics by solving the ODE:
@@ 4.2 Two-stage training scheme
Supervised fine-tuning. We first conduct a supervised fine-tuning (SFT) on our ConStoryBoard dataset to pre-train the STEP2 model, establishing a strong foundation for inter-shot transitions. Specifically, we employ the Low-Rank Adaptation (LoRA) technique [lora] on both the understanding model $\mathcal{E}_{mu}$ (for high-level semantic understanding) and the generation model $\mathcal{E}_{gen}$ (for low-level frame generation). Following the flow matching formulation [liu2022flow], our training target is the constant velocity vector field $v_{t}=x_{i}^{1}-x_{i}^{0}$, pointing from the noise $x_{i}^{0}\sim\mathcal{N}(0,\mathbf{I})$ to the ground-truth latent pair $x_{i}^{1}$:
@@
Preference alignment. To further align our STEP2 model with human preferences, we perform the post-training using Direct Preference Optimization (DPO) [dpo]. Using the SFT-trained model as the reference model $v_{\text{ref}}$ and extracting preference tuple $(y_{w},y_{l})$ from the ConStoryBoard-HP subset, this process aims to maximize the likelihood of policy model $v_{\theta}$ prefers $y_{w}$ over $y_{l}$:
@@ 4.3 STAGE workflow
1) User’s Inputs and Storyboard Generation: Given a user-provided text-based story theme $T_{\text{desc}}$, we design a Director Agent $G_{\text{dire}}$ with chain-of-thought [cot] pre-defined prompts to generate a structured storyboard $\mathcal{S}$:
@@
Subsequently, this refined description $R_{i}$ and the start-end frame pair $(F^{\mathrm{S}}_{i},F^{\mathrm{E}}_{i})$ are fed into an off-the-shelf video generation model $G_{\text{video}}$ (e.g., WanX [wanx], Veo3.1 [veo3]) to guide the generation process of the video clip $V_{i}$ of the $i$-th shot:

### v2

@@ Abstract
While recent advancements in generative models have achieved remarkable visual fidelity in video synthesis, creating coherent multi-shot narratives remains a significant challenge. To address this, keyframe-based approaches have emerged as a promising alternative to computationally intensive end-to-end methods, offering the advantages of fine-grained control and greater efficiency. However, these methods often fail to maintain cross-shot consistency and capture cinematic language. In this paper, we introduce STAGE, a SToryboard-Anchored GEneration workflow to reformulate the keyframe-based multi-shot video generation task. Instead of using sparse keyframes, we propose STEP2 to predict a structural storyboard composed of start-end frame pairs for each shot. We introduce the multi-shot memory pack to ensure long-range entity consistency, the dual-encoding strategy for intra-shot coherence, and the two-stage training scheme to learn cinematic inter-shot transition. We also contribute the large-scale ConStoryBoard dataset, including high-quality movie clips with fine-grained annotations for story progression, cinematic attributes, and human preferences. Extensive experiments demonstrate that STAGE achieves superior performance in structured narrative control and cross-shot coherence. Our code will be available at this url.
Figure 1: Illustration of our proposed STAGE workflow for multi-shot video generation. Given a user-provided story description, existing end-to-end [52] and keyframe-based [13] methods often suffer from incoherent inter-shot transitions (e.g., motion discontinuities and appearance inconsistencies) that disrupt the narrative flow. We address this by explicitly predicting a structural storyboard composed of start-end frame pairs, which serve as visual anchors to ensure superior long-range shot consistency.
@@ 1 Introduction
Recent advancements in generative models have dramatically improved the visual quality of synthesized videos, fueling their adoption on short-clip content platforms [40, 39, 17]. To further unlock their potential for more sophisticated applications (e.g., filmmaking), recent models are evolving to enhance storytelling ability [60, 56, 27]. However, this narrative ability typically requires composing a long video from multiple distinct shots depicting different stages or perspectives of a story, which still poses significant challenges for existing models.
While directly generating an entire multi-shot video in an end-to-end manner [52, 46, 15, 9] is an intuitive approach, it is computationally expensive and runs in an “all-or-nothing” paradigm, leading to an inefficient trial-and-error process with limited user control. To provide fine-grained user control, keyframe-based approaches [8, 51, 59, 58] first generate several keyframes to establish the video’s narrative structure, and subsequently employ external I2V models [54, 35, 3] to synthesize each shot accordingly. However, these methods suffer from disrupted cross-shot coherence and struggle to capture cinematic language (Fig.˜1 (a), the close-up reveals motion discontinuity when opening the ring box).
@@
To address this challenge, we introduce STAGE, a workflow for SToryboard-Anchored GEneration for cinematic multi-shot narrative. The technical core of this workflow is our STart-End frame-Pair Prediction model (STEP2), which iteratively visualizes the text-based storyboard into $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$ pairs. To further enable STEP2 to visualize the storyboard of each shot effectively, we propose: (i) a multi-shot memory pack to compress the history and ensure long-range entity consistency; (ii) a dual-encoding strategy to enforce intra-shot coherence and logical correlation; and (iii) a two-stage training scheme to understand the complex cinematic language of inter-shot transitions. In the complete STAGE workflow, a director agent first augments the user-provided story theme into a text-based storyboard. Our fully-trained STEP2 model then iteratively generates the $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$ pairs. These pairs are finally fed into an off-the-shelf video generation model (e.g., WanX [45], Veo3.1 [43]) to generate video clips, which are then concatenated into the final multi-shot video.
To facilitate the training of STEP2, we collect open-sourced movie clips [2] and construct the ConStoryBoard dataset [^1] We will publish the dataset once the paper is accepted.. It consists of 100K video clips with start-end frame pairs, along with structural storyboard annotations (e.g., story progress of each shot) and diverse cinematic language attributes (e.g., shot scale, shot length, camera angle, and camera movement). Building on this, we further curate the ConStoryBoard-HP by manually selecting high-quality frame pairs and constructing preference tuples for human-preference alignment post-training.
@@ 2.1 Single-Shot video generation
The field of single-shot video generation has seen remarkable progress. Building on the success of image diffusion models [34, 36, 31, 4], early methods [10, 38, 11] adapt pre-trained models by incorporating temporal layers. However, these approaches often result in limited temporal coherence. To address these limitations, research has shifted towards Diffusion Transformer architectures [28]. While state-of-the-art video foundation models [57, 18, 45] are able to generate high-fidelity videos with a single shot, they typically lack a native mechanism for multi-shot narratives. Consequently, naively applying them shot-by-shot results in jarring visual discontinuities that disrupt the narrative flow, marking the key challenge our work addresses.
@@ 2.2 Multi-Shot video generation
Generating multi-shot videos requires maintaining both content consistency and narrative coherence across shots. Existing methods extend single-shot models to generate full sequences, either via end-to-end visual conditioning [24, 58] or by modifying attention mechanisms [15, 52, 25]. These approaches are computationally expensive and offer limited user control. To address these limitations, an alternative approach [60, 56, 51, 30] first generates sparse keyframes and then synthesizes the full video clips based on them. Despite their advances in text fidelity and story coherence, they largely neglect the cinematic language for inter-shot transitions, leading to abrupt cuts and logical disconnects that break the narrative flow. In contrast, our STAGE explicitly reformulates the keyframe-based approach as a start-end frame pair prediction problem, thereby directly modeling these inter-shot transitions.
@@ 2.3 Reinforcement learning for visual generation
Inspired by the success of Reinforcement Learning from Human Feedback (RLHF) in language models, aligning visual models with human preferences is a promising research area. Initial efforts involve directly fine-tuning models with scalar reward signals [55, 5] or employing reward weighted regression [29, 20]. Subsequently, policy gradient algorithms (e.g., PPO [37]) are integrated into diffusion models, demonstrating notable efficacy in visual quality improvement [7, 26]. To avoid significant computational overhead and training instability, Direct Preference Optimization (DPO) [33] emerges and demonstrates its effectiveness in aligning visual models with human preferences for static image properties, such as text-image fidelity [44, 47] and aesthetic quality [21]. Building on this, we extend this technique to optimize for the complex temporal relationships (e.g., cinematic language) that define coherent storytelling in multi-shot videos.
@@ 3 Dataset
Current keyframe-based video datasets [52, 53, 30] offer a single keyframe for each shot, primarily focusing on text-to-shot alignment while neglecting the start and end frames for modeling inter-shot transitions. This motivates us to construct the ConStoryBoard dataset, tailored to train our start-end frame pair prediction model.
The construction of ConStoryBoard begins with multi-shot videos collected from the Condensed Movies dataset [2], which we filter for high-resolution (over 1080p) and high aesthetic scores [19] (over 5.5). These filtered videos are then segmented into individual video clips for each shot using TransNetV2 [41]. For each resulting clip ($i$-th shot in the original video), we employ InternVL-3.5 [48] to generate a detailed text description $D_{i}$ of the story progression, and structured cinematic attributes $C_{i}$ (e.g., shot scale and camera movement). We then extract the ground-truth start-end frame pair $(F^{\mathrm{S}}_{i},F^{\mathrm{E}}_{i})$ from these annotated clips, and post-process them by cropping black bars and removing watermarks. Finally, the ConStoryBoard dataset comprises 100K training pairs and 1K testing pairs, where each sample consists of a ground-truth pair $(F^{\mathrm{S}}_{i},F^{\mathrm{E}}_{i})$ and its corresponding annotations $(D_{i},C_{i})$.
@@ 4.1 Start-end frame pair prediction model
Specifically, as illustrated in Fig.˜3 (a), given preceding shot start-end frame pairs $\{(F_{j}^{\mathrm{S}},F_{j}^{\mathrm{E}})\}_{j=1}^{i-1}$ for $i$-th shot, we disorderly collect all frames in these pairs into a memory bank $\{F_{j}^{\mathrm{M}}\}_{j=1}^{2i-2}$, and encode all these frames into the latent space $\{m_{j}\}_{j=1}^{2i-2}=\{\mathcal{E}_{\mathrm{vae}}(F_{j}^{\mathrm{M}})\}_{j=1}^{2i-2}$ using a pre-trained VAE encoder $\mathcal{E}_{\mathrm{vae}}$. To guide the iterative generation with semantically relevant memory cues, we rank these preceding latent codes based on their CLIP similarity [32] as $\{m^{\prime}_{j}\}_{j=1}^{2i-2}=\mathrm{Sort}\big(\{m_{j}\}_{j=1}^{2i-2}\big)$. After that, this ranked memory bank is further compressed into a packed memory token via a progressive spatial tiling mechanism:
@@
Specifically, as shown in Fig.˜3 (b), we encode the ground truth start-end frame pair $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$ separately using a pre-trained VAE encoder $\mathcal{E}_{vae}$, and then concatenate them along the sequence dimension to construct a single joint shot tensor $x_{i}=\big[\mathcal{E}_{\mathrm{vae}}(F_{i}^{\mathrm{S}});\mathcal{E}_{\mathrm{vae}}(F_{i}^{\mathrm{E}})\big]$. As the common practice of flow matching [22], this shot tensor is then linearly interpolated with a Gaussian noise as:
@@
Multi-modal understanding and generation. To reason about shot semantics from diverse context, our STEP2 first adopts a multi-modal large language model to understand the story progress and character performance of each shot. Specifically, as illustrated in Fig.˜3 (c), the understanding model $\mathcal{E}_{mu}$ is built upon Qwen2.5-VL [1], which leverages the end frame of the previous shot $F_{i-1}^{\mathrm{E}}$, the $i$-th shot corresponding text description $D_{i}$, and cinematic attributes $C_{i}$ to generate unified context tokens $U_{i}=\mathcal{E}_{mu}(F_{i-1}^{\mathrm{E}},D_{i},C_{i})$ for the $i$-th shot.
With the aforementioned memory tokens $M_{i}$ and interpolated joint shot tensor $x_{i}^{t}$, these unified context tokens $U_{i}$ are fed into the generation model $\mathcal{E}_{gen}$ to generate the corresponding start-end frame pair $(F_{i}^{\mathrm{S}},F_{i}^{\mathrm{E}})$. Specifically, the generation model $\mathcal{E}_{gen}$ is implemented as a diffusion Transformer architecture, comprising multiple MMDiT blocks [6] that includes the self-attention layers for global context interaction. This enables our STEP2 to generate clean start-end frame pairs referring all integrated and compression semantics by solving the ODE:
@@ 4.2 Two-stage training scheme
Supervised fine-tuning. We first conduct a supervised fine-tuning (SFT) on our ConStoryBoard dataset to pre-train the STEP2 model, establishing a strong foundation for inter-shot transitions. Specifically, we employ the Low-Rank Adaptation (LoRA) technique [12] on both the understanding model $\mathcal{E}_{mu}$ (for high-level semantic understanding) and the generation model $\mathcal{E}_{gen}$ (for low-level frame generation). Following the flow matching formulation [23], our training target is the constant velocity vector field $v_{t}=x_{i}^{1}-x_{i}^{0}$, pointing from the noise $x_{i}^{0}\sim\mathcal{N}(0,\mathbf{I})$ to the ground-truth latent pair $x_{i}^{1}$:
@@
Preference alignment. To further align our STEP2 model with human preferences, we perform the post-training using Direct Preference Optimization (DPO) [33]. Using the SFT-trained model as the reference model $v_{\text{ref}}$ and extracting preference tuple $(y_{w},y_{l})$ from the ConStoryBoard-HP subset, this process aims to maximize the likelihood of policy model $v_{\theta}$ prefers $y_{w}$ over $y_{l}$:
@@ 4.3 STAGE workflow
1) User’s Inputs and Storyboard Generation: Given a user-provided text-based story theme $T_{\text{desc}}$, we design a Director Agent $G_{\text{dire}}$ with chain-of-thought [49] pre-defined prompts to generate a structured storyboard $\mathcal{S}$:
@@
Subsequently, this refined description $R_{i}$ and the start-end frame pair $(F^{\mathrm{S}}_{i},F^{\mathrm{E}}_{i})$ are fed into an off-the-shelf video generation model $G_{\text{video}}$ (e.g., WanX [45], Veo3.1 [43]) to guide the generation process of the video clip $V_{i}$ of the $i$-th shot: